consider the reaction
KMnO4 + 5FeCl2 + 8HCl =MnCl2 + 5FeCl3 + 4H2O + KCl
if 24.0 mol of FeCl2 and 13.0 mol KMnO4 are mixed with excess HCl, how many moles of MnCl2 can be formed
1
Expert's answer
2021-08-28T06:27:21-0400

Mole ratio of MnCl2 :KMnO4 = 1:1

13 KMnO4 = 13MnCl2

Mole ratio of MnCl2 : FeCl2 = 1:5

= 24/5 = 4.8 moles

Mole ratio of MnCl2: HCl = 1:8

= 1/8 = 0.125

= 13+0.125 + 4.8

= 17.925 moles